# Balancing Intensity and Focality in Directional DBS Under Uncertainty: A Simulation Study of Electrode Optimization via a Metaheuristic L1L1 Approach

### v1

@@ 2.5 Uncertainty Modeling
Uncertainty refers to various physiological and technical factors that directly or indirectly affect the accuracy and reliability of the stimulation, including imprecision in lead placement, patient-specific anatomical variation, and inhomogeneities in tissue conductivity. In this study, all sources of uncertainty affecting the optimization problem are modeled as additive perturbations to the lead field. While this Gaussian noise model is a simplification, it enables a tractable first-step analysis of robustness. We acknowledge that in practice, the true uncertainty arises from a variety of unrelated sources.
@@
Figure 2: (2(a)): A schematic of how the peak signal-to-noise ratio (PSNR) $\delta^{-1}$ relates to the set $\mathcal{V}_{\delta}$ and to the approximate maximum dynamic range $\varepsilon^{-1}$ of the volumetric stimulation current field. (2(b)): The set $\mathcal{V}_{\delta}$ shown for PSNR 10–40 dB, providing a rough approximation of the maximum feasible coverage of volume of tissue activated (VTA) for an activation threshold set by $\delta$.
8-contact Electrode Configuration
40-contact Electrode Configuration
Figure 3: Examples of current injection patterns obtained using the Reciprocity Principle (RP), metaheuristic L1-norm regularized L1-norm fitting (L1L1) method under two different nuisance current thresholds (L1L1(A) with $\varepsilon=\{0:-160\}$ and L1L1(B) with $\varepsilon=\{0:-10\}$), and the Tikhonov-Regularized Least Square (TLS) method. The target position and orientation (parallel to the lead) is indicated by the red upward pin. The bar plots display the resulting electrode current amplitudes, while the contact maps visualize their spatial distribution over the electrodes; cathodal (blue), anodal (red), and inactive (gray/green). The RP method yields a bipolar configuration maximizing focused current density $\Gamma$, while the remaining methods searches for a multipolar configuration maximizing focality $\Theta=\Gamma/\Xi$ under an a priori constraint $\varepsilon$ on the nuisance current density $\Xi$, determined by the peak signal-to-noise ratio (PSNR) $\delta^{-1}$ and the lead field attenuation. The L1L1(B) result reflects more strongly constrained $\Xi$ (smaller dynamic range), activating fewer contacts whereas L1L1(A) and TLS allow higher $\Theta$ values and broader excitation patterns. The voltage/current on the electrode contacts are ±$\pm 2.0$ mA (milliamperes).
We postulate that the volumetric current density distribution can be optimized for those degrees of freedom indexed by $i$, where the L2-norm (2-norm) $\|{\bf\tilde{l}}_{i}\|_{2}$ of the corresponding reduced lead field matrix row remains above the threshold. This row, denoted as ${\bf l}_{i}$, is first reduced (as described in Appendix 9) and then zero-averaged, taking the form
where $\overline{L}_{i}=K^{-1}\sum_{j=1}^{K}L_{i,j}$ represents the mean of the row entries. The threshold for optimization is determined by a peak signal-to-noise ratio (PSNR) constraint, ensuring that only degrees of freedom with a sufficiently strong contribution to the lead field are considered in the optimization process. We define the feasible set of degrees of freedom as
where $\max_{k}\|{\bf\tilde{l}}_{k}\|_{2}$ is the maximum L2-norm of the zero-averaged lead field rows. The parameter $\delta$ sets a relative threshold, filtering out degrees of freedom whose contributions are too weak to be considered meaningful. If $\delta$ is chosen to represent neural activation threshold, then $\mathcal{V}_{\delta}$ serves as an approximation of the maximum VTA (Fig. 2(a)).
Since lead field attenuation and uncertainty impose fundamental limits on stimulation accuracy, the PSNR constraint defines the theoretical limit of the optimization performance. If the target stimulation site corresponds to the $\ell$-th index, then the maximum achievable dynamic range $\varepsilon^{-1}$ between the current density at the target site and its surroundings is approximately
This relationship indicates that the minimum contrast in current density between the target location and surrounding regions is constrained by $\delta$ and the structure of the lead field matrix (Fig. 2(b)).
@@ 2.6 Numerical Experiments
Our numerical experiments are designed to evaluate the performance of the proposed optimization methods with respect to: (i) simulated 8-contact and 40-contact DBS lead configurations, (ii) noiseless and noisy forward mappings, (iii) orientation of the current dipole in the model (parallel or perpendicular with respect to the lead array) [17], and (iv) algorithm efficiency. The hardware used for these experiments includes a Dell Precision 5820 Workstation with 256 GB of RAM, a 10-core Intel i9–10900X CPU, and an NVidia Quadro RTX 4000 GPU. The GPU (graphics processing unit) was utilized to accelerate both forward and inverse computations.
Table 1: Summary of the proposed optimization methods. While the RP provides a bipolar reciprocal of the target, both L1L1 and TLS constitute a metaheuristic search over a two-dimensional hyperparameter lattice. L1L1 is subdivided into L1L1(A) and L1L1(B) based on two alternative a priori assumptions of the feasible $\varepsilon$ threshold range which is [-160, 0] and [-10, 0] dB, respectively. The regularization and the L2-norm weighting parameter intervals are chosen as in [32]. CPU time refers to the computing time needed to evaluate a single optimizer with fixed hyperparameter values for HR-LF with the 40-contact lead configuration.
@@
The L1L1 method is examined using two alternative a priori assumptions for the feasible $\varepsilon$ threshold range: [-160, 0] dB and [-10, 0] dB, respectively. We refer to these settings as ”L1L1(A)” and ”L1L1(B)”. The former setting neglects the effect of uncertainty on the field ratio $\Theta$, while the second assumes that uncertainty is included within 10 dB dynamic range. As for TLS method, the regularization parameter $\gamma$ and the L2-norm weighting parameter $\beta$ intervals are chosen as in [32] (Table 1).
Gaussian noise is added to the lead field to examine spatial sensitivity; the effect of noise is assessed in an overall setting, where the candidate solution is determined for each target DOF of LR-LF under PSNR 40.0 dB noise corruption. While anatomical variations may cause spatially correlated errors, the implementation of noise offers a generic and tractable way to assess noise-robustness when the exact structure of uncertainty is unknown. It serves as a standard baseline for evaluating sensitivity to model perturbations.
The noise sensitivity is examined for two specific target positions (I) and (II) (Fig. 4) using PSNRs of 40.0 and 50.0 dB, along the HR-LF. When interpreted in terms of $\mathcal{V}_{\delta}$ (Figure 2(b)), PSNR 40.0 dB (50.0 dB) corresponds to the assumption that each possible target position in the set $\mathcal{V}_{-30\text{ dB}}$ ($\mathcal{V}_{-40\text{ dB}}$) can be stimulated with at least a 10 dB dynamic range $\varepsilon^{-1}$ for $\Theta$. This assumption aligns well with the $\varepsilon$ range of $[-10,0]$ selected for L1L1(B).
Figure 4: Visualization of the two synthetic stimulation targets, Target (I) and Target (II), analyzed separately in numerical experiments. Both targets are positioned within the right thalamus near a directional DBS lead. The figure also displays the lead’s center point and a concentric spherical region of interest (ROI) with 6 mm radius, representing an approximate VTA. This ROI defines the region where maximizing stimulation intensity while minimizing current spread to surrounding areas is considered feasible. As targets (I) and (II) are not right next to the probe, lead field uncertainty affects the optimization process. Bilateral thalamic structures are shown in gray for anatomical reference.
@@
Figure 5: Volumetric optimization results for the 8-contact probe and noiseless lead fields. Decision variable $\Gamma$, $\Xi$ and $\Theta$ for optimizers found via RP, TLS, L1L1(A), and L1L1(B) method illustrated as a function of the target position and orientation (parallel or perpendicular w.r.t. the probe). Left: distributions for a (i) Low-Resolution Lead Field (LR-LF) where the spatial degrees of freedom (DOFs) cover both thalamus lobes. Right: distribution for a (ii) High-Resolution Lead Field (HR-LF) where the DOFs is confined to a region of interest within a radius of 6.0 mm (millimeters) from the center of the lead. The optimal electrode configuration found by the L1L1 method maximizes the field ratio $\Theta$ subject to $\Gamma>=\Gamma_{0}$. The $\Theta$ distribution found using L1L1(A) is elevated as compared to the case of RP and TLS. The limited dynamic range ($\varepsilon$ range) of L1L1(B) is reflected in an overall slightly lower $\Theta$.
@@ 3 Results
The applied algorithms generate distinct distributions of the variables $\Gamma$, $\Xi$, and $\Theta$, which can be differentiated based on how they are spatially organized across the target domain under the optimal electrode configurations for the 8- (Fig. 5) and 40-contact (Fig. 6) leads. The direction of the dipole current plays a key role in shaping electric field targeting. The anode–cathode pair exhibiting the strongest positive and negative values may activate electrodes located as close as $0.189\pm 0.009\,\text{mm}$ or as far as $6.375\pm 0.375\,\text{mm}$ apart across the lead. This effect becomes evident when comparing the corresponding distributions: in the parallel orientation, the targeting is highly concentrated near the tip of the probe, whereas the perpendicular orientation results in a broader spatial spread, with activity shifted medially relative to the lead.
@@
The bipolar configurations obtained via RP yield the overall highest focused current density $\Gamma$; however, this comes at the expense of focality, as indicated by elevated nuisance current $\Xi$ and a reduced field ratio $\Theta$. This trade-off is understandable, as the multipolar configurations are designed to maximize the field ratio $\Theta$ subject to the constraint $\Gamma\geq\Gamma_{0}>0$, with $\Gamma_{0}=0.8\,\text{mA}$, thereby allowing activation of more than two contacts. In the noiseless setting, the L1L1(B) reconstructions with $\varepsilon\in[-10,0]\,\text{dB}$ yield slightly higher values of $\Gamma$ and $\Xi$, but exhibit a more moderate $\Theta$ compared to L1L1(A) with $\varepsilon\in[-160,0]\,\text{dB}$. Nonetheless, both L1L1(A) and L1L1(B) demonstrate superior nuisance current suppression relative to TLS and RP.
Figure 7 and Table 2 compare the noiseless and noisy ($\text{PSNR}=40\,\text{dB}$) results for $\Gamma$ and $\Theta$ obtained using L1L1(A) and L1L1(B) and LR-LF. The results suggest that, unlike L1L1(A), L1L1(B) is more robust to lead field noise, yielding similar outcomes in both settings. In contrast, L1L1(A), which uses a smaller lower bound for $\varepsilon$, tends to overfit the noiseless data. This leads to unrealistically high values of the field ratio $\Theta$: up to $27.07$ in the 8-contact configuration and as high as $536.05$ in the 40-contact case. However, in the presence of noise, the results produced by L1L1(A) fall within a similar range to those of L1L1(B), indicating that the overfitting is specific to the noiseless scenario.
@@
Figure 8 further examines the noise sensitivity of RP, TLS, L1L1(A), and L1L1(B) for two specific target positions, illustrated in Figure 4, covering a sample of twenty different noise realizations for both PSNR 40 and 50 dB. The results demonstrate that in a noisy situation, L1L1(A) and L1L1(B) provide similar distributions for $\Gamma$ and $\Theta$, hence, also for $\Xi$. Compared to RP and TLS, the L1L1 approach seems to provide a somewhat elevated field ratio $\Theta$ with a difference that is only marginal or absent with PSNR 40 dB, but observable with PSNR 50 dB. Accordingly, the potential benefit of using L1L1 depends critically on the level of modeling uncertainty, and it vanishes when this uncertainty becomes excessive. Of the four optimization approaches tested, L1L1(B) provides the closest agreement between noiseless and noisy outcomes and proves particularly effective in maximizing the field ratio $\Theta$ in the presence of noise.
@@
Table 2: The maximum values of the decision variable (DV) $\Theta$ (field ratio) found by L1L1(A) and L1L1(B) searches and the corresponding values of $\Gamma$ (focused current density) for the low-resolution lead field (LR-LF) and its target set (Figure 7) covering both parallel and perpendicular target orientations of the DBS lead in both noiseless and noisy (PSNR 40 dB) setting.
@@
  Noiseless lead field Refer to caption A1. Focus $\Gamma$ (A/m2)     Refer to caption B1. Ratio $\Theta$ (rel.) Noisy lead field Refer to caption C1. Focus $\Gamma$ (A/m2)     Refer to caption D1. Ratio $\Theta$ (rel.)
@@
  Noiseless lead field Refer to caption A2. Focus $\Gamma$ (A/m2)     Refer to caption B2. Ratio $\Theta$ (rel.) Noisy lead field Refer to caption C2. Focus $\Gamma$ (A/m2)     Refer to caption D2. Ratio $\Theta$ (rel.)
@@
  Parallel Refer to caption Focus $\Gamma$ (A/m2) Refer to caption Ratio $\Theta$ (rel.) Perpendicular Refer to caption Focus $\Gamma$ (A/m2) Refer to caption Ratio $\Theta$ (rel.) Target (I) Parallel Refer to caption Focus $\Gamma$ (A/m2) Refer to caption Ratio $\Theta$ (rel.) Perpendicular Refer to caption Focus $\Gamma$ (A/m2) Refer to caption Ratio $\Theta$ (rel.) Target (II)
@@
  Parallel Refer to caption Focus $\Gamma$ (A/m2) Refer to caption Ratio $\Theta$ (rel.) Perpendicular Refer to caption Focus $\Gamma$ (A/m2) Refer to caption Ratio $\Theta$ (rel.) Target (I) Parallel Refer to caption Focus $\Gamma$ (A/m2) Refer to caption Ratio $\Theta$ (rel.) Perpendicular Refer to caption Focus $\Gamma$ (A/m2) Refer to caption Ratio $\Theta$ (rel.) Target (II)
Figure 8: Boxplots showing the focused current density $\Gamma$ and field ratio $\Theta$ obtained with a noise-corrupted HR-LF and the 8-contact electrode configuration for test targets (I) and (II). The optimizers have been found through RP, TLS, L1L1(A), and L1L1(B) search. Each bar collates a sample of twenty decision variable values corresponding to optimizers obtained with as many independent lead field noise realizations (PSNR 40 or 50 dB). The box shows the interquartile range (IQR) between the first quantile $Q1$ and the third one $Q3$, with a horizontal line at the median. The whiskers extend from the edges of the box to the smallest and largest values within $[Q1-1.5\times\text{IQR},Q3+1.5\times\text{IQR}]$, where $\text{IQR}=Q3-Q1$. Data points outside this range are considered outliers and are plotted individually as markers. The results obtained with L1L1(A) and L1L1(B) demonstrate a comparable sensitivity under lead field noise, suggesting that the limited dynamic range of L1L1(B) does not downgrade the L1L1 method’s performance in situations involving uncertainty. Furthermore, both L1L1(A) and L1L1(B) find an overall elevated field ratio compared to RP and TLS with the greater PSNR level 50 dB, which allows for an improved +10 dB dynamic range for the optimization algorithm as compared to PSNR 40 dB.

### v2

@@ 2.4 Uncertainty Modeling
Uncertainty refers to various physiological and technical factors that affect the accuracy and reliability of the stimulation, including imprecision in lead placement, patient-specific anatomical variation, and inhomogeneities in tissue conductivity [schmidt2016uncertainty, athawale2019statistical]. Potential sources affecting the optimization framework are modeled as additive perturbations to the lead field matrix through Gaussian noise model. While this is a simplification, it enables a tractable first-step analysis of robustness. We acknowledge that in practice, the true uncertainty arises from a variety of unrelated sources.
We postulate that the volumetric current density distribution can be optimized for DOFs indexed by $i$ whose corresponding reduced lead field row ${\bf\tilde{l}}_{i}$ has an L2-norm above a threshold. This row is first reduced and zero-averaged, taking the form
where $\overline{L}_{i}=K^{-1}\sum_{j=1}^{K}L_{i,j}$ is the mean of the row entries. The threshold is determined via a peak signal-to-noise ratio (PSNR) constraint, ensuring that only DOFs contributing significantly to the lead field are included in the optimization. The feasible set of DOFs is defined as
where $\max_{k}\|{\bf\tilde{l}}_{k}\|_{2}$ is the largest L2-norm among all the zero-averaged rows in the lead field rows. The parameter $\delta$ sets a relative threshold, filtering out DOFs whose contributions are too weak to be meaningful. When $\delta$ is interpreted as the activation threshold, $\mathcal{V}_{\delta}$ approximates the maximum VTA (Fig. 3(a)).
Because lead field attenuation and uncertainty impose limits on stimulation accuracy, the PSNR constraint effectively defines the theoretical performance limit of the optimization. If the target site corresponds to index $\ell$, the maximum achievable dynamic range $\varepsilon^{-1}$ between the current density at the target and surrounding regions is approximately
This indicates that the minimum contrast between the target and non-target regions is constrained by both $\delta$ and the structure of the lead field matrix (Fig. 3(b)).
@@
Figure 3: (3(a)): A schematic of how the peak signal-to-noise ratio (PSNR) $\delta^{-1}$ relates to the set $\mathcal{V}_{\delta}$ and to the approximate maximum dynamic range $\varepsilon^{-1}$ of the volumetric stimulation current field. (3(b)): The set $\mathcal{V}_{\delta}$ shown for PSNR 10–40 dB, providing a rough approximation of the maximum feasible coverage of volume of tissue activated (VTA) for an activation threshold set by $\delta$.
2.5 Hardware
Dell Precision 5820 Workstation equipped with 256 GB of RAM, a 10-core Intel i9–10900X CPU, and an NVidia Quadro RTX 4000 GPU. The GPU were used to accelerate both forward and inverse calculations.
@@ 2.6 Numerical Experiments
Our numerical experiments evaluate the proposed optimization methods with respect to: (i) 8- and 40-contact DBS lead configurations, (ii) noiseless and noisy forward mappings, (iii) dipole orientation relative to the lead array (parallel or perpendicular) [Janson_2020_Activation], and (iv) algorithm efficiency.
Table 1: Summary of the proposed optimization methods. RP provides a bipolar reciprocal of the target, whereas L1L1 and TLS perform an extensive metaheuristic search. L1L1 method is subdivided into L1L1(A) and L1L1(B) according to two alternative a priori assumptions for the feasible $\varepsilon$ threshold range: [-160, 0] dB and [-10, 0] dB, respectively. The regularization and L2-norm weighting parameter ranges are adopted from [galazprieto_2022_L1vsL2]. CPU time refers to the computation time needed to evaluate a single optimizer with fixed hyperparameters for the HR-LF using the 40-contact lead.
@@
The L1L1 method is tested under two alternative a priori assumptions for the feasible $\varepsilon$ threshold range: [-160, 0] dB (L1L1(A)) and [-10, 0] dB (L1L1(B)). The first neglects uncertainty in the field ratio $\Theta$, while the second accounts for variability within a 10 dB dynamic range. TLS uses regularization parameter $\gamma$ and L2-norm weighting parameter $\beta$ intervals as in [galazprieto_2022_L1vsL2] (Table 1). Detailed formulation of both the L1L1 and TLS methods is provided in A.2 and A.3, respectively.
To examine sensitivity to measurement and modeling errors, Gaussian noise is added to the lead field. Candidate solutions are computed for each target DOF of the LR-LF under PSNR 40.0 dB corruption. While anatomical variations can introduce spatially correlated errors, this additive noise provides a tractable baseline to assess robustness to model perturbations. Noise sensitivity is specifically analyzed for two target positions (I) and (II) (Fig. 4) along the HR-LF using PSNRs of 40.0 and 50.0 dB. Interpreted in terms of $\mathcal{V}_{\delta}$, PSNR 40.0 dB (50.0 dB) corresponds to the assumption that each target in $\mathcal{V}_{-30\text{ dB}}$ ($\mathcal{V}_{-40\text{ dB}}$) can be stimulated with at least a 10 dB dynamic range $\varepsilon^{-1}$ for $\Theta$, which aligns with the $\varepsilon$ range of [-10, 0] used for L1L1(B).
Figure 4: Visualization of the two synthetic stimulation targets, Target (I) and Target (II), analyzed separately in numerical experiments. Both targets are positioned within the right thalamus near a directional DBS lead. The figure also displays the lead’s center point and a concentric spherical region of interest (ROI) with 6 mm radius, representing an approximate VTA. This ROI defines the region where maximizing stimulation intensity while minimizing current spread to surrounding areas is considered feasible. As targets (I) and (II) are not right next to the probe, lead field uncertainty affects the optimization process. Bilateral thalamic structures are shown in gray for anatomical reference. The distances are shown with respect to the boundary cylindrical probe illustrated by the ring around the center point.
@@
Figure 5: Volumetric optimization results for the 8-contact probe and noiseless lead fields. Decision variable $\Gamma$, $\Xi$ and $\Theta$ for optimizers found via Reciprocity Principle (RP), L1-norm regularized L1-norm fitting (L1L1) evaluated as L1L1(A) and L1L1(B) with feasible $\varepsilon$ threshold range: [-160, 0] dB and [-10, 0] dB, respectively, and Tikhonov Regularized Least Squares (TLS) methods illustrated as a function of the target position and orientation (parallel or perpendicular w.r.t. the probe). Left: distributions for a Low-Resolution Lead Field (LR-LF) where the spatial degrees of freedom (DOFs) cover both thalamus lobes. Right: distribution for a High-Resolution Lead Field (HR-LF) where the DOFs is confined to a region of interest within a radius of 6.0 mm (millimeters) from the center of the lead. The optimal electrode configuration found by the L1L1 method maximizes the field ratio $\Theta$ subject to $\Gamma>=\Gamma_{0}$. The $\Theta$ distribution found using L1L1(A) is elevated as compared to the case of RP and TLS. The limited dynamic range ($\varepsilon$ range) of L1L1(B) is reflected in an overall slightly lower $\Theta$.
@@ 3 Results
The applied algorithms generate distinct distributions of the variables $\Gamma$, $\Xi$, and $\Theta$, which can be differentiated based on how they are spatially organized across the target domain under the optimal electrode configurations for the 8- (Fig. 5) and 40-contact (Fig. 6) leads. The direction of the current dipole plays the role in shaping electric field targeting. The anode–cathode pair exhibiting the strongest positive and negative values may activate electrodes located as close as $0.189\pm 0.009\,\text{mm}$ or as far as $6.375\pm 0.375\,\text{mm}$ apart across the lead. With the parallel orientation, the targeting is highly concentrated near the tip of the probe, whereas the perpendicular orientation results in a broader spatial spread, with activity shifted medially relative to the lead.
@@
Bipolar configurations obtained via RP yield the highest focused current density $\Gamma$; however, this comes at the expense of focality, as indicated by elevated nuisance current $\Xi$ and a reduced field ratio $\Theta$. This trade-off is understandable, as the multipolar configurations are designed to maximize the field ratio $\Theta$ subject to the constraint $\Gamma\geq\Gamma_{0}>0$, with $\Gamma_{0}=0.8\,\text{mA}$, thereby allowing activation of more than two contacts.
Figure 7 and Table 2 compare noiseless and noisy ($\text{PSNR}=40\,\text{dB}$) results for $\Gamma$ and $\Theta$ using the L1L1 method over a LR-LF. In the noiseless setting, the L1L1(B) reconstructions with $\varepsilon\in[-10,0]\,\text{dB}$ yield slightly higher $\Gamma$ and $\Xi$ values, and moderate $\Theta$ compared to L1L1(A) with $\varepsilon\in[-160,0]\,\text{dB}$. L1L1(A) provided unrealistically high values of the field ratio $\Theta$: up to $27.07$ in the 8-contact directional DBS leads and as high as $536.05$ in the 40-contact case indicating a case of overfitting.
@@
Figure 8 further examines the noise sensitivity of the applied optimization methods using two specific target positions, illustrated in Figure 4, covering a sample of twenty different noise realizations for both PSNR 40 and 50 dB. The results demonstrate that in a noisy situation, L1L1(A) and L1L1(B) provide similar distributions for $\Gamma$ and $\Theta$. Compared to RP and TLS, the L1L1 approach seems to provide elevated field ratio $\Theta$ with a difference that is only marginal or absent with PSNR 40 dB, but observable with PSNR 50 dB. Accordingly, the potential benefit of using L1L1 depends critically on the level of modeling uncertainty, and it vanishes when this uncertainty becomes excessive. Of the four optimization approaches tested, L1L1(B) provides the closest agreement between noiseless and noisy outcomes and proves particularly effective in maximizing the field ratio $\Theta$ in the presence of noise.
@@
Table 2: Maximum focused current density and field ratio described as $\Gamma_{max}$ and $\Theta_{max}$, respectively, for a 8- and 40-contacts multipolar DBS lead using the L1L1 method over a low-resolution lead field (LR-LF), with current dipole targeting aligned parallel and perpendicular to the DBS lead, for both noiseless (No Noise) and noisy (PSNR 40 dB) settings.
@@
Noiseless lead field
A1. Focus $\Gamma$ (A/m2)
B1. Ratio $\Theta$ (rel.)
Noisy lead field
C1. Focus $\Gamma$ (A/m2)
D1. Ratio $\Theta$ (rel.)
@@
Noiseless lead field
A2. Focus $\Gamma$ (A/m2)
B2. Ratio $\Theta$ (rel.)
Noisy lead field
C2. Focus $\Gamma$ (A/m2)
D2. Ratio $\Theta$ (rel.)
@@
Parallel
Focus $\Gamma$ (A/m2)
Ratio $\Theta$ (rel.)
Perpendicular
Focus $\Gamma$ (A/m2)
Ratio $\Theta$ (rel.)
Target (I)
Parallel
Focus $\Gamma$ (A/m2)
Ratio $\Theta$ (rel.)
Perpendicular
Focus $\Gamma$ (A/m2)
Ratio $\Theta$ (rel.)
Target (II)
@@
Parallel
Focus $\Gamma$ (A/m2)
Ratio $\Theta$ (rel.)
Perpendicular
Focus $\Gamma$ (A/m2)
Ratio $\Theta$ (rel.)
Target (I)
Parallel
Focus $\Gamma$ (A/m2)
Ratio $\Theta$ (rel.)
Perpendicular
Focus $\Gamma$ (A/m2)
Ratio $\Theta$ (rel.)
Target (II)
Figure 8: Boxplots showing the focused current density $\Gamma$ and field ratio $\Theta$ obtained with a noise-corrupted HR-LF and the 8-contact electrode configuration for test targets (I) and (II). The optimizers have been found through RP, L1L1(A), L1L1(B), TLS methods. Each bar collates a sample of twenty decision variable values corresponding to optimizers obtained with as many independent lead field noise realizations (PSNR 40 or 50 dB). The box shows the interquartile range (IQR) between the first quantile $Q1$ and the third one $Q3$, with a horizontal line at the median. The whiskers extend from the edges of the box to the smallest and largest values within $Q1-1.5\times\text{IQR},Q3+1.5\times\text{IQR}]$, where $\text{IQR}=Q3-Q1$. Data points outside this range are considered outliers and are plotted individually as markers. The results obtained with L1L1(A) and L1L1(B) demonstrate a comparable sensitivity under lead field noise, suggesting that the limited dynamic range of L1L1(B) does not downgrade the L1L1 method’s performance in situations involving uncertainty. Furthermore, both L1L1(A) and L1L1(B) find an overall elevated field ratio compared to RP and TLS with the greater PSNR level 50 dB, which allows for an improved +10 dB dynamic range for the optimization algorithm as compared to PSNR 40 dB.